# Performance of Quantum Approximate Optimization with Quantum Error Detection

### v1

@@ Abstract
Quantum algorithms must be scaled up to tackle real-world applications. Doing so requires overcoming the noise present on today’s hardware. The quantum approximate optimization algorithm (QAOA) is a promising candidate for scaling up due to its modest resource requirements and documented asymptotic speedup over state-of-the-art classical algorithms for some problems. However, achieving better-than-classical performance with QAOA is believed to require fault tolerance. In this paper, we demonstrate a partially fault-tolerant implementation of QAOA using the $[[k+2,k,2]]$ “Iceberg” error detection code. We observe that encoding the circuit with the Iceberg code improves the algorithmic performance as compared to the unencoded circuit for problems with up to $20$ logical qubits on a trapped-ion quantum computer. Additionally, we propose and calibrate a model for predicting the code performance, and use it to characterize the limits of the Iceberg code and extrapolate its performance to future hardware with improved error rates. In particular, we show how our model can be used to determine necessary conditions for QAOA to outperform Goemans-Williamson algorithm on future hardware. Our results demonstrate the largest universal quantum computing algorithm protected by partially fault-tolerant quantum error detection on practical applications to date, paving the way towards solving real-world applications with quantum computers.
@@ I Introduction
Quantum computers are poised to deliver algorithmic speedups for a broad range of application in science and industry [1, 2, 3]. However, realizing these speedups requires overcoming the challenge presented by the noise which limits the computational power of today’s quantum devices. Error correction [4] provides a scalable path to fault-tolerance and has shown significant progress in hardware recently [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Nonetheless, quantum error-correction imposes large overheads, making it challenging to execute even small-scale applications fully fault-tolerantly. As a result, fully fault-tolerant demonstrations of quantum algorithms for practical applications have been out of reach of experiment, despite immense progress in implementation and benchmarking of algorithmic components, such as preparation of magic states [15], one-bit addition [16] and quantum Fourier transform [17].
@@
Quantum error detection (QED) codes provide an opportunity for partially fault-tolerant implementation of algorithms in the near term [9, 10, 18, 19, 20, 21, 22]. While non-scalable, they can still deliver improved algorithmic performance beyond what is possible without protecting against noise [23, 24, 25, 26]. The protection offered by QED codes opens an opportunity to use quantum computers to study the performance of quantum algorithms for sizes and noise rates beyond classical simulation.
@@
The recently proposed $[[k+2,k,2]]$ “Iceberg” QED code [18] is particularly suitable for near-term algorithms due to its ability to encode expressive circuits, using a universal set of local and global logical rotations, with a low overhead. The Iceberg code has been demonstrated to improve the fidelity of random circuits with up to $8$ logical qubits and $1323$ physical two-qubit gates [18], the performance of quantum phase estimation with up to $4$ logical qubits and $920$ two-qubit gates [25], and the fidelity of ground state preparation with probabilistic imaginary-time evolution with $4$ logical qubits and up to $906$ two-qubit gates [26]. While not fully fault-tolerant, these experiments provided preliminary evidence that for circuits with small numbers of qubits the Iceberg code can improve algorithmic performance.
@@
Quantum Approximate Optimization Algorithm (QAOA) [27, 28] is a quantum optimization heuristic applicable to a broad range of combinatorial optimization problems in finance and other industries [29, 30, 31, 32]. QAOA has been shown to provide a quantum algorithmic speedup over state-of-the-art solvers for some problems [33, 34], motivating its implementation on hardware. While relatively low resource requirements enabled QAOA execution on non-error-corrected hardware [35, 36, 37, 38, 39, 40, 41, 42], realizing the speedup offered by QAOA is widely believed to require fault tolerance [43, 44]. We remark that both quantum hardware performance and the impact of quantum noise on QAOA have been subject of extensive interest [45, 46, 47, 48, 49, 50, 51, 52].
@@
We demonstrate a partially fault-tolerant implementation of QAOA applied to the MaxCut problem on Quantinuum H2-1 trapped-ion quantum computer [54] with the Iceberg code. At the time of experiments, the H2-1 device had $32$ all-to-all connected qubits and 99.8% two-qubit gate fidelity [38]. We execute circuits with up to $24$ logical qubits encoded into up to $26$ physical qubits using up to $813$ physical two-qubit gates and observe that protecting the circuit with the Iceberg code leads to improved approximation ratio as compared to the unencoded circuit for problems requiring up to $20$ logical qubits. However, we also observe that beyond $20$ logical qubits, Iceberg code does not yield an improved algorithmic performance. To the best of our knowledge, these experiments are the largest evaluation of Iceberg code and the largest QED-encoded application demonstration in terms of the number of logical qubits to date.
To understand the protection capability of the Iceberg code, we propose a model that predicts the code performance. The model efficiently constructs an analytical estimation of the logical fidelity and post-selection rate as a function of the circuit size and three error rates related to the noise produced by two-qubit physical gates. We calibrate the model by simulating a large set of QAOA circuits of varying sizes with and without the Iceberg code in the emulator [55, 56, 8] of the H2-1 quantum computer. The calibrated model is then used it to characterize the regimes in which the Iceberg code improves the algorithmic performance of QAOA. Specifically, we identify the ranges for the number of logical qubits, QAOA depth, and the number of syndrome measurements for which Iceberg code is beneficial. Furthermore, our model can be used to predict the performance of the Iceberg code on future improved hardware. We demonstrate conditions on effective error rates for QAOA to outperform the Goemans-Williamson (GW) [57] algorithm on small graphs.
@@ II Background
Given a graph $G(k,E)$ with $k$ vertices and set of edges $E$, the MaxCut problem consists in finding a cut that partitions the vertices into two sets that maximise the number of edges between them. Cuts can be represented by strings $\bm{z}$ of $k$ bits with value $z_{i}=\pm 1$ if vertex $i$ is in one set or the other. The MaxCut objective function can be written as $f(z)=\sum_{(i,j)\in E}(1-z_{i}z_{j}$).
@@ II.1 Quantum Approximate Optimization Algorithm
QAOA is a quantum algorithm for combinatorial optimization. It solves optimization problems by preparing quantum state using a sequence of $\ell$ layers of alternating cost Hamiltonian and mixing Hamiltonian operators, parameterized by vectors $\bm{\gamma}$ and $\bm{\beta}$ respectively.
@@
The parameters $\bm{\gamma}$ and $\bm{\beta}$ are chosen such that the measurement outcomes of |$\ket{\bm{\psi}}$ correspond to high-quality solutions of the optimization problem with high probability. In this paper, we take the initial state |$\ket{\bm{\psi}_{0}}=\ket{+}^{\otimes k}$ as the equal superposition of all possible candidate solutions, and the mixing Hamiltonian as a summation of all single-qubit Pauli-$X$ operator $\mathcal{M}=\sum_{i=1}^{n}X_{i}$.
@@
Recent progress in parameter setting heuristics has considerably advanced the execution of QAOA in the early fault-tolerant era [58, 42], with good parameter choices available for many problems. A set of parameters that leads to good approximation ratios was proposed in [59] for the MaxCut problems on regular graphs that we solve in this work. Throughout our paper we use these “fixed angles” to set QAOA parameters in the experiments.
@@ II.2 Iceberg Code
To detect these errors, the fault-tolerant initialization, syndrome measurement and final measurement blocks, employ two ancillas. In the absence of noise, the state remains purely in the $+1$ eigenspace of the stabilizers during the entire circuit execution and ancillas always output a $+1$ when measured. The final measurement block additionally measures the stabilizer $S_{Z}$, which is also expected to be measured as $+1$ in the absence of noise. Therefore, a $-1$ output in any of them signals the presence of an error caused by noise, and the circuit execution is discarded.
The fault-tolerant design of the initialization, syndrome measurement and final measurement blocks ensures that no single faulty component in these blocks (like a two-qubit gate) can cause a logical error. In contrast, our logical gates, despite being natural for the hardware, are not fault-tolerant, as some errors in their physical implementation can not be detected. We nevertheless show in Sec. III.2.2 that undetectable errors are rare, rendering the QAOA protection of the Iceberg code effectively fully fault-tolerant.
@@ III.1 Iceberg code protection of QAOA on hardware
The performance of the Iceberg code with QAOA on 3-regular graph MaxCut on the Quantinuum H2-1 quantum computer [54] is shown in Fig. 1C. The logical fidelity reported in this figure is estimated from the average energy measured experimentally by assuming a global white noise model distribution, as described in Sec. V.3. We fix the QAOA depth to $\ell=10$ and vary the number of logical qubits $k$, randomly selecting one MaxCut graph instance per $k$. For each problem we run QAOA unencoded and QAOA protected by the Iceberg code with three intermediate syndrome measurements. Throughout this paper, the final measurement is counted as a syndrome measurement, so in the previous experiment we say that four syndrome measurements are used. The Iceberg data has larger error bars due to the smaller number of post-selected samples.
@@
The histogram in Fig. 1D reports the hardware shots of several Iceberg and unencoded QAOA circuits for $k=16$. After post-selection, the output distribution has higher weight on bitstrings with higher approximation ratio, as expected from a better protection against noise.
@@
We compare the performance with that of the Pauli check sandwiching (PCS) [53, 60], an error detection scheme with a similar motivation to that of the Iceberg code. PCS uses pairs of parity checks to detect some but not necessarily all errors that occur in a given part of the circuit. The parity checks are chosen based on the symmetries already present in the circuit. For QAOA circuits considered in this work, the problem Hamiltonian commutes with $X^{\otimes k}$ and $Z^{\otimes k}$, so we use them as the checks of our PCS experiments. To unify the notation with Iceberg code, a pair of $X^{\otimes k}$ and $Z^{\otimes k}$ checks is denoted as one syndrome measurement. For example, one syndrome measurement in PCS means that we select one cost Hamiltonian layer $e^{-i\gamma\mathcal{H}}$ and sandwich it with two parity checks. The overhead of one syndrome measurement in PCS includes an additional $4k$ two-qubit gates, along with two ancillas. The comparison with PCS on a $k=18$, $\ell=11$ QAOA circuit is shown in Fig. 1B and E, where all data are from the H2-1 emulator (H2-1E). We observe that PCS leads to a lower logical fidelity that does not increase with the number of syndrome measurements. At this scale, the large overhead and the non-fault tolerant design of the PCS method decreases the circuit performance. At the same time, we observe that the Iceberg code can effectively improve the QAOA performance and obtain a higher logical fidelity than the unencoded circuit with four syndrome measurements.
@@ III.2 Estimated performance from our model
To understand the protection capability of Iceberg code, we propose the performance model of Sec. V.2 for the unencoded and Iceberg code circuits. The model outputs analytical functions of the logical fidelity $\mathcal{F}_{\mathrm{une}}$ for the unencoded circuits, the logical fidelity $\mathcal{F}_{\mathrm{ice}}$ for the Iceberg code circuits, and the post-selection rate $1-D$ for the Iceberg code ($D$ is the discard rate). Inputs from the circuit are the numbers of logical qubits $k$, logical single-qubit gates $g_{1}$, logical two-qubit gates $g_{2}$, and syndrome measurements $s$. From the hardware, the model for the unencoded and the Iceberg code respectively inputs only one and three error rates related to the noise produced by two-qubit physical gates. These are motivated and described in more detail in the next section. In this work we leave the error rates as fitting parameters so that, when fitted to data from the H2-1 emulator the fitted values incorporate corrections from other noise sources.
@@ III.2.1 Dataset and model validation
From this dataset we select partial data that have relatively stable logical fidelity and large numbers of two-qubit gates to fit the model. For the Iceberg code model, that has three fitting parameters, we use data from $64$ encoded circuits while the data from $15$ unencoded circuits is used to fit the unencoded model, that has only one fitting parameter. We additionally filter out those Hamiltonian terms of every QAOA circuit whose expected values are outliers with respect to the white noise approximation. More details are provided in Appendix B. The mean and $95\%$ confidence interval of the fitted parameters from $3000$ boostrapping iterations are reported in Table 1.
Figure 3: Model prediction: Predicting the performance of QAOA with the number of logical qubits in the range $k\in[6,48]$ and the number of QAOA layers in the range $\ell\in[1,16]$. We use the model proposed in this work to estimate (A, B) the difference $\mathcal{F}_{\mathrm{ice}}-\mathcal{F}_{\mathrm{une}}$ in logical fidelity between the Iceberg and the unencoded circuits, and to estimate (C, D) the post-selection rate. In A and C, we use the model fitted error rates in Table 1 and vary the number of syndrome measurements. In B and D, we fix the number of syndrome measurements at 4 and scale down the model error rates by the indicated factors. The red lines in the top row (A, B) show where the logical fidelity of Iceberg code circuits equals that of unencoded circuits. The orange lines in the bottom row (C, D) indicate where the Iceberg code circuits have a 10% post-selection rate.
To validate the accuracy of the performance model, we present the logical fidelity and post-selection rate of experimental data alongside the model predictions in Fig. 2. We find that our model can match the experimental results both qualitatively and quantitatively. The model and experiment fidelities and post-selection rates for every selected circuit, as well as the deviations from the white noise simplification are presented in Appendix B.

### v3

@@ Abstract
Quantum algorithms must be scaled up to tackle real-world applications. Doing so requires overcoming the noise present on today’s hardware. The quantum approximate optimization algorithm (QAOA) is a promising candidate for scaling up, due to its modest resource requirements and documented asymptotic speedup over state-of-the-art classical algorithms for some problems. However, achieving better-than-classical performance with QAOA is believed to require fault tolerance. In this paper, we demonstrate a partially fault-tolerant implementation of QAOA using the $[[k+2,k,2]]$ “Iceberg” error detection code. We observe that encoding the circuit with the Iceberg code improves the algorithmic performance as compared to the unencoded circuit for problems with up to $20$ logical qubits on a trapped-ion quantum computer. Additionally, we propose and calibrate a model for predicting the code performance. We use this model to characterize the limits of the Iceberg code and extrapolate its performance to future hardware with improved error rates. In particular, we show how our model can be used to determine the necessary conditions for QAOA to outperform the Goemans-Williamson algorithm on future hardware. To the best of our knowledge, our results demonstrate the largest universal quantum computing algorithm protected by partially fault-tolerant quantum error detection on practical applications to date, paving the way towards solving real-world applications with quantum computers.
@@ I Introduction
Quantum computers are poised to deliver algorithmic speedups for a broad range of applications in science and industry [1, 2, 3, 4]. However, realizing these speedups requires overcoming the challenge presented by the noise which limits the computational power of today’s quantum devices. Error correction [5] provides a scalable path to fault-tolerance and has shown significant progress in hardware recently [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Nonetheless, quantum error-correction imposes large overheads, making it challenging to execute even small-scale applications fully fault-tolerantly. As a result, fully fault-tolerant demonstrations of quantum algorithms for practical applications have been out of reach of experiments, despite the immense progress in implementation and benchmarking of algorithmic components, such as preparation of magic states [16], one-bit addition [17], and quantum Fourier transform [18].
@@
Quantum error-detection (QED) codes provide an opportunity for partially fault-tolerant implementation of algorithms in the near-term [10, 11, 19, 20, 21, 22, 23]. While non-scalable, they can still deliver improved algorithmic performance beyond what is possible without protecting against noise [24, 25, 26, 27]. The protection offered by QED codes opens an opportunity to use quantum computers to study the performance of quantum algorithms for sizes and noise rates beyond classical simulation.
@@
The recently proposed $[[k+2,k,2]]$ “Iceberg” QED code [19] is particularly suitable for near-term algorithms due to its ability to encode expressive circuits, using a universal set of local and global logical rotations, with a low overhead. The Iceberg code has been demonstrated to improve the fidelity of random circuits with up to $8$ logical qubits and $1,323$ physical two-qubit gates [19], the performance of quantum phase estimation with up to $4$ logical qubits and $920$ two-qubit gates [26], and the fidelity of ground state preparation with probabilistic imaginary-time evolution with $4$ logical qubits and up to $906$ two-qubit gates [27]. While not fully fault-tolerant, these experiments provided preliminary evidence that for circuits with small numbers of qubits, the Iceberg code can improve algorithmic performance.
@@
Quantum Approximate Optimization Algorithm (QAOA) [28, 29] is a quantum optimization heuristic applicable to a broad range of combinatorial optimization problems in finance and other industries [30, 31, 32, 33]. QAOA has been shown to provide a quantum algorithmic speedup over state-of-the-art solvers for some problems [34, 35], motivating its implementation on hardware. While relatively low resource requirements enabled QAOA execution on non-error-corrected hardware [36, 37, 38, 39, 40, 41, 42, 43, 44], realizing the speedup offered by QAOA is widely believed to require fault tolerance [45, 46, 47]. We remark that both quantum hardware performance and the impact of quantum noise on QAOA have been subjects of extensive interest [48, 49, 50, 51, 52, 53, 54, 55].
@@
We demonstrate a partially fault-tolerant implementation of QAOA applied to the MaxCut problem on the Quantinuum H2-1 trapped-ion quantum computer [57] with the Iceberg code. At the time of our experiments, the H2-1 device had $32$ all-to-all connected qubits and 99.8% two-qubit gate fidelity [39]. We execute circuits with up to $24$ logical qubits encoded into up to $26$ physical qubits using up to $813$ physical two-qubit gates. With this, we observe that protecting the circuit with the Iceberg code leads to an improved approximation ratio, as compared to the unencoded circuit for problems requiring up to $20$ logical qubits. However, we also observe that beyond $20$ logical qubits, the Iceberg code does not yield an improved algorithmic performance. To the best of our knowledge, these experiments are the largest evaluation of the Iceberg code and the largest QED-encoded application demonstration in terms of the number of logical qubits to date.
To understand the protection capability of the Iceberg code, we propose a model that predicts the code performance. The model efficiently constructs an analytical estimation of the logical fidelity and post-selection rate as a function of the circuit size and three error rates related to the noise produced by two-qubit physical gates. We calibrate the model by simulating a large set of QAOA circuits of varying sizes, with and without the Iceberg code in the emulator [58, 59, 9] of the H2-1 quantum computer. The calibrated model is then used to characterize the regimes in which the Iceberg code improves the algorithmic performance of QAOA. Specifically, we identify the ranges for the number of logical qubits, QAOA depth, and the number of syndrome measurements for which the Iceberg code is beneficial. Furthermore, our model can be used to predict the performance of the Iceberg code on future improved hardware. We demonstrate conditions on effective error rates for QAOA to outperform the Goemans-Williamson (GW) [60] algorithm on small graphs.
@@ II Background
Given a graph $G(k,E)$ with $k$ vertices and set of edges $E$, the MaxCut problem consists in finding a cut that partitions the vertices into two sets that maximize the number of edges between them. Cuts can be represented by strings $\bm{z}$ of $k$ bits with value $z_{i}=\pm 1$ if vertex $i$ is in one set or the other. The MaxCut objective function can be written as $f(z)=\sum_{(i,j)\in E}(1-z_{i}z_{j}$).
@@ II.1 Quantum Approximate Optimization Algorithm
QAOA is a quantum algorithm for combinatorial optimization. It solves optimization problems by preparing a quantum state using a sequence of $\ell$ layers of alternating cost Hamiltonian and mixing Hamiltonian operators, parameterized by vectors $\bm{\gamma}$ and $\bm{\beta}$, respectively.
@@
The parameters $\bm{\gamma}$ and $\bm{\beta}$ are chosen such that the measurement outcomes of |$\ket{\bm{\psi}}$ correspond to high-quality solutions of the optimization problem with high probability. In this paper, we take the initial state |$\ket{\bm{\psi}_{0}}=\ket{+}^{\otimes k}$ as the equal superposition of all possible candidate solutions, and the mixing Hamiltonian as a summation of all single-qubit Pauli-$X$ operators $\mathcal{M}=\sum_{i=1}^{n}X_{i}$.
@@
Recent progress in parameter setting heuristics has considerably advanced the execution of QAOA in the early fault-tolerant era [61, 43], with good parameter choices available for many problems. A set of parameters that leads to good approximation ratios was proposed in [62] for the MaxCut problems on regular graphs that we solve in this work. Throughout our paper, we use these “fixed angles” to set QAOA parameters in the experiments.
@@ II.2 Iceberg Code
To detect these errors, the fault-tolerant initialization, syndrome measurement, and final measurement blocks employ two ancillas. In the absence of noise, the state remains purely in the $+1$ eigenspace of the stabilizers during the entire circuit execution and ancillas always output a $+1$ when measured. The final measurement block additionally measures the stabilizer $S_{Z}$, which is also expected to be measured as $+1$ in the absence of noise. Therefore, a $-1$ output in any of them signals the presence of an error caused by noise, and the circuit execution is discarded.
The fault-tolerant design of the initialization, syndrome measurement, and final measurement blocks ensures that no single faulty component in these blocks (like a two-qubit gate) can cause a logical error. In contrast, our logical gates, despite being natural for the hardware, are not fault-tolerant, as some errors in their physical implementation cannot be detected. Nevertheless, we show in Sec. III.2.2 that undetectable errors are rare, rendering the QAOA protection of the Iceberg code effectively fully fault-tolerant.
@@ III.1 Iceberg code protection of QAOA on hardware
The performance of the Iceberg code with QAOA on 3-regular graph MaxCut on the Quantinuum H2-1 quantum computer [57] is shown in Fig. 1C. The logical fidelity reported in this figure is estimated from the average energy measured experimentally by assuming a global white noise model distribution, as described in Sec. V.3. We fix the QAOA depth to $\ell=10$ and vary the number of logical qubits $k$, randomly selecting one MaxCut graph instance per $k$. For each problem we run QAOA unencoded and QAOA protected by the Iceberg code with three intermediate syndrome measurements. Throughout this paper, the final measurement is counted as a syndrome measurement, so in the previous experiment we say that four syndrome measurements are used. The Iceberg data has larger error bars due to the smaller number of post-selected samples. The performance of Iceberg circuits decreases as $k$ increases. Meanwhile, the performance of unencoded circuits is relatively robust because the accumulated error of additional gates when increasing $k$ is not significant to diminish performance at this scale.
@@
The histogram in Fig. 1D reports the hardware shots of several Iceberg and unencoded QAOA circuits for $k=16$. After post-selection, the output distribution has higher weight on bitstrings with a higher approximation ratio, as expected from a better protection against noise.
@@
We compare the performance with that of the Pauli check sandwiching (PCS) [56, 63], an error detection scheme with a similar motivation to that of the Iceberg code. PCS uses pairs of parity checks to detect some but not necessarily all errors that occur in a given part of the circuit. The parity checks are chosen based on the symmetries already present in the circuit. For QAOA circuits considered in this work, the problem Hamiltonian commutes with $X^{\otimes k}$ and $Z^{\otimes k}$, so we use them as the checks of our PCS experiments. To unify the notation with Iceberg code, a pair of $X^{\otimes k}$ and $Z^{\otimes k}$ checks is denoted as one syndrome measurement. For example, one syndrome measurement in PCS means that we select one cost Hamiltonian layer $e^{-i\gamma\mathcal{H}}$ and sandwich it with two parity checks. The overhead of one syndrome measurement in PCS includes an additional $4k$ two-qubit gates, along with two ancillas. The comparison with PCS on a $k=18$, $\ell=11$ QAOA circuit is shown in Fig. 1B and E, where all data are from the H2-1 emulator (H2-1E). We observe that PCS leads to a lower logical fidelity that does not increase with the number of syndrome measurements. At this scale, the large overhead and the non-fault tolerant design of the PCS method decreases the circuit performance. At the same time, we observe that the Iceberg code can effectively improve the QAOA performance and obtain a higher logical fidelity than the unencoded circuit with four syndrome measurements. Here, we do not aim to claim QED is better than quantum error mitigation techniques as QEM usually aims to improve the accuracy of estimating an observable, rather than improving the quality of a single bitstring outcome. This study does not aim to claim the superiority of QED over quantum error mitigation (QEM) techniques, as QEM is generally designed to enhance the accuracy of observable estimations rather than the quality of individual bitstring outcomes. PCS was chosen as the baseline due to its unique capability among QEM methods to perform error detection, which aligns with the principles of Iceberg encoding, albeit without the necessity of encoding.
@@ III.2 Estimated performance from our model
To understand the protection capability of Iceberg code, we propose the performance model of Sec. V.2 for the unencoded and Iceberg code circuits. The model outputs analytical functions of the logical fidelity $\mathcal{F}_{\mathrm{une}}$ for the unencoded circuits, the logical fidelity $\mathcal{F}_{\mathrm{ice}}$ for the Iceberg code circuits, and the post-selection rate $1-D$ for the Iceberg code ($D$ is the discard rate). Inputs from the circuit are the numbers of logical qubits $k$, logical single-qubit gates $g_{1}$, logical two-qubit gates $g_{2}$, and syndrome measurements $s$. From the hardware, the model for the unencoded and the Iceberg code respectively inputs only one and three error rates related to the noise produced by two-qubit physical gates. These are motivated and described in more detail in the next section. We leave the error rates as fitting parameters so that, when fitted to data from the H2-1 emulator, the fitted values incorporate corrections from other noise sources.
@@ III.2.1 Dataset and model validation
From this dataset we select partial data that have relatively stable logical fidelity and large numbers of two-qubit gates to fit the model. For the Iceberg code model, which has three fitting parameters, we use data from $64$ encoded circuits. The data from $15$ unencoded circuits is used to fit the unencoded model, which has only one fitting parameter. We additionally filter out those Hamiltonian terms of every QAOA circuit whose expected values are outliers with respect to the white noise approximation. More details are provided in Appendix B. The mean and $95\%$ confidence interval of the fitted parameters from $3000$ boostrapping iterations are reported in Table 1.
Figure 3: Model prediction: Predicting the performance of QAOA with the number of logical qubits in the range $k\in[6,48]$ and the number of QAOA layers in the range $\ell\in[1,16]$. We use the model proposed in this work to estimate (A, B) the difference $\mathcal{F}_{\mathrm{ice}}-\mathcal{F}_{\mathrm{une}}$ in logical fidelity between the Iceberg and the unencoded circuits, and to estimate (C, D) the post-selection rate. In A and C, we use the model fitted error rates in Table 1 and vary the number of syndrome measurements. In B and D, we fix the number of syndrome measurements at 4 and scale down the model error rates by the indicated factors. The red lines in the top row (A, B) show where the logical fidelity of Iceberg code circuits equals that of unencoded circuits. The cyan lines in the bottom row (C, D) indicate where the Iceberg code circuits have a 10% post-selection rate.
To validate the accuracy of the performance model, we present the logical fidelity and post-selection rate of emulated data alongside the model predictions in Fig. 2. We find that our model can match the emulated results both qualitatively and quantitatively. The model and experimental fidelities and post-selection rates for every selected circuit, as well as the deviations from the white noise simplification, are presented in Appendix B.